# Beyond Gaps and Bumps: Spectral Siren Cosmology with Non-Parametric Population Models

### v1

@@ Abstract
Gravitational wave standard sirens typically require electromagnetic (EM) data to obtain redshift information to constrain cosmology. Difficult to find EM counterparts for bright sirens and galaxy survey systematics for dark sirens make cosmological constraints with spectral sirens, a gravitational wave data-only approach, extremely appealing. In this work, we use the GWTC-3 BBH detections as spectral sirens to constrain the BBH population and the underlying cosmological expansion with a flexible model for the black hole mass spectrum. We use a binned Gaussian process to model the BBH mass distribution in the source frame without any astrophysical assumptions on the shape and or inclusion (or lack of) features that drive the cosmological constraints as the redshifted detector frame masses become consistent with the underlying astrophysical mass distribution features. For GWTC-3 we find a measurement on the Hubble constant of $H_{0}=73.0^{+13.3}_{-7.7}$ $\mbox{km s}^{-1}\,\mbox{Mpc}^{-1}$ at $68\%$ C.L. when combined with that obtained from the bright standard siren analysis with GW170817 and its associated host galaxy NGC 4993. We find an improved estimate for the Hubble constant of around a factor of 1.4 times better than the GW170817 measurement alone. We validate our nonparametric spectral siren approach with simulations and benchmark its scalability and constraining performance when compared with parametric methods. $H_{0}=73.0^{+13.3}_{-7.7}\ {\rm{km\ s^{-1}\ Mpc^{-1}}}$
@@ 1 Introduction
In this work, we employ, for the first time, a flexible population model on the BBH population distribution and perform hierarchical Bayesian inference on the GWTC-3 BBH detections to measure the cosmic expansion history using the spectral siren methodology. We model the BBH population using the binned gaussian process (BGP) approach (Ray et al., 2023; Mohite, 2022; Abbott et al., 2023a) to place constraints on the Hubble constant without assuming a particular shape on the mass distribution of BBH mergers. For a discussion regarding how our method compares with a parallel investigation on non-parametric spectral sirens (Farah et al., In prep.), see Sec. 5.
@@ 5 Discussion and Conclusion
In a parallel investigation, Farah et al. (In prep.) have also attempted to use Gaussian-process-based mass models on a simulated dataset for a proof of concept demonstration of astrophysics agnostic spectral siren cosmology. Even though their non-parametric population model is not restricted by bin resolutions, the mock data analysis presented in Farah et al. (In prep.) models only the primary mass distribution of CBCs with gaussian processes. They further assume that all BBHs are equal mass systems and use a parametric model for the redshift evolution of the merger rate. By simulating a large population of equal mass BBHs with a known functional form of the redshift evolution of the merger rate, they demonstrate that their method can implement accurate spectral siren cosmology in a simplistic Universe while being free of the necessity to bin up the space of BBH parameters. We note that while an unbinned population model in both component masses and redshifts could in principle be more flexible than ours, Farah et al. (In prep.) have not demonstrated whether their method can be scalably generalized and applied to more realistic datasets wherein BBH masses are expected to follow more complicated pairing functions (Abbott et al., 2023a) than a delta function in mass ratio centered around one. Furthermore, given the strong correlations between the measurement uncertainties of the Hubble parameter and the redshift evolution parameters (Abbott et al., 2021e), previously unmodelled features in the redshift evolution function can serve to equally bias the cosmological parameters. On the other hand, our model can reconstruct features in the joint distribution of component masses and the redshift evolution function completely driven by the data, up to the resolution limit imposed by our choice of binning, variations of which are found to yield consistent results.
@@ Appendix A Convergence of Monte-Carlo integrals
$\sum_{\gamma,\alpha}\left(n^{\gamma}_{m}n^{\alpha}_{z}\sigma_{w,i}^{\gamma\alpha}(\Omega)\right)^{2}$ $=$ $\left\langle\left(\frac{\frac{dN}{dm_{1}dm_{2}dz}(\vec{n}_{m},\vec{n_{z}})}{p_{\text{PE}}(m_{1},m_{2},z)}\right)^{2}\right\rangle_{\text{samples},i}-\frac{1}{N_{samples,i}}\left(\sum_{\gamma,\alpha}n^{\gamma}_{m}n^{\alpha}_{z}\mu_{w,i}^{\gamma\alpha}(\Omega)\right)^{2}$ (A2)
@@
$\frac{N_{eff}^{\rm{VT}}(\Omega)}{N_{det}(\vec{n}_{m},\vec{n_{z}},\Omega)}=\frac{\sum_{\gamma,\alpha}\left(\frac{\mu_{\rm{VT}}^{\gamma\alpha}}{\sigma_{\rm{VT}}^{\gamma\alpha}}\right)^{2}}{\sum_{\gamma,\alpha}n^{\gamma}_{m}n^{\alpha}_{z}\mu_{\rm{VT}}^{\gamma\alpha}}$$\geq$ 22 22 (A5)
$\min_{i}N_{eff}^{w,i}(\vec{n}_{m},\vec{n}_{z},\Omega)=\min_{i}\left\{\frac{\left(\sum_{\gamma,\alpha}n^{\gamma}_{m}n^{\alpha}_{z}\mu_{w,i}^{\gamma\alpha}\right)^{2}}{\sum_{\gamma,\alpha}(n^{\gamma}_{m}n^{\alpha}_{z}\sigma_{w,i}^{\gamma\alpha})^{2}}\right\}$$\geq$ $10^{0.6}$ (A6)

### v2

@@ Abstract
Gravitational wave standard sirens typically require electromagnetic (EM) data to obtain redshift information to constrain cosmology. Difficult to find EM counterparts for bright sirens and galaxy survey systematics for dark sirens make cosmological constraints with spectral sirens, a gravitational wave data-only approach, extremely appealing. In this work, we use the GWTC-3 BBH detections as spectral sirens to constrain the BBH population and the underlying cosmological expansion with a flexible model for the black hole mass spectrum. We use a binned Gaussian process to model the BBH mass distribution in the source frame without any astrophysical assumptions on the shape and or inclusion (or lack of) features that drive the cosmological constraints as the redshifted detector frame masses become consistent with the underlying astrophysical mass distribution features. For GWTC-3 we find a measurement on the Hubble constant of $H_{0}=73.0^{+13.3}_{-7.7}$ $\mbox{km s}^{-1}\,\mbox{Mpc}^{-1}$ at $68\%$ C.L. when combined with that obtained from the bright standard siren analysis with GW170817 and its associated host galaxy NGC 4993. We find an improved estimate for the Hubble constant of around a factor of 1.4 times better than the GW170817 measurement alone. We validate our nonparametric spectral siren approach with simulations and benchmark its scalability and constraining performance when compared with parametric methods.
@@ 1 Introduction
In this work, we employ, for the first time, a flexible population model on the BBH population distribution and perform hierarchical Bayesian inference on the GWTC-3 BBH detections to measure the cosmic expansion history using the spectral siren methodology. We model the BBH population using the binned gaussian process (BGP) approach (Ray et al., 2023, 2024; Mohite, 2022; Abbott et al., 2023a) to place constraints on the Hubble constant without assuming a particular shape on the mass distribution of BBH mergers. For a discussion regarding how our method compares with a parallel investigation on non-parametric spectral sirens (Farah et al., 2024), see Sec. 5.
@@ 5 Discussion and Conclusion
In a parallel investigation, Farah et al. (2024) have also attempted to use Gaussian-process-based mass models on a simulated dataset for a proof of concept demonstration of astrophysics agnostic spectral siren cosmology. Even though their non-parametric population model is not restricted by bin resolutions, the mock data analysis presented in Farah et al. (2024) models only the primary mass distribution of CBCs with gaussian processes. They further assume that all BBHs are equal mass systems and use a parametric model for the redshift evolution of the merger rate. By simulating a large population of equal mass BBHs with a known functional form of the redshift evolution of the merger rate, they demonstrate that their method can implement accurate spectral siren cosmology in a simplistic Universe while being free of the necessity to bin up the space of BBH parameters. We note that while an unbinned population model in both component masses and redshifts could in principle be more flexible than ours, Farah et al. (2024) have not demonstrated whether their method can be scalably generalized and applied to more realistic datasets wherein BBH masses are expected to follow more complicated pairing functions (Abbott et al., 2023a) than a delta function in mass ratio centered around one. Furthermore, given the strong correlations between the measurement uncertainties of the Hubble parameter and the redshift evolution parameters (Abbott et al., 2021e), previously unmodelled features in the redshift evolution function can serve to equally bias the cosmological parameters. On the other hand, our model can reconstruct features in the joint distribution of component masses and the redshift evolution function completely driven by the data, up to the resolution limit imposed by our choice of binning, variations of which are found to yield consistent results.
@@ Appendix A Convergence of Monte-Carlo integrals
$\sum_{\gamma,\alpha}\left(n^{\gamma}_{m}n^{\alpha}_{z}\sigma_{w,i}^{\gamma\alpha}(\Omega)\right)^{2}$ $=$ $\left\langle\left(\frac{\frac{dN}{dm_{1}dm_{2}dz}(\vec{n}_{m},\vec{n_{z}})}{p_{\text{PE}}(m_{1},m_{2},z)}\right)^{2}\right\rangle_{\text{samples},i}-\frac{1}{N_{\text{samples},i}}\left(\sum_{\gamma,\alpha}n^{\gamma}_{m}n^{\alpha}_{z}\mu_{w,i}^{\gamma\alpha}(\Omega)\right)^{2}$ (A2)
@@
$\frac{N_{\rm{eff}}^{\rm{VT}}(\Omega)}{N_{\rm{det}}(\vec{n}_{m},\vec{n_{z}},\Omega)}=\frac{\sum_{\gamma,\alpha}\left(\frac{\mu_{\rm{VT}}^{\gamma\alpha}}{\sigma_{\rm{VT}}^{\gamma\alpha}}\right)^{2}}{\sum_{\gamma,\alpha}n^{\gamma}_{m}n^{\alpha}_{z}\mu_{\rm{VT}}^{\gamma\alpha}}$$\geq$ 22 22 (A5)
$\min_{i}N_{\rm{eff}}^{w,i}(\vec{n}_{m},\vec{n}_{z},\Omega)=\min_{i}\left\{\frac{\left(\sum_{\gamma,\alpha}n^{\gamma}_{m}n^{\alpha}_{z}\mu_{w,i}^{\gamma\alpha}\right)^{2}}{\sum_{\gamma,\alpha}(n^{\gamma}_{m}n^{\alpha}_{z}\sigma_{w,i}^{\gamma\alpha})^{2}}\right\}$$\geq$ $10^{0.6}$ (A6)